# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@ 5.1 Shimura variety attached to a lattice
Let $\mathbf{D}=\mathbf{D}(M)$ be the type IV Hermitian symmetric domain associated to $M$ defined by
@@
$\mathbf{D}(M)=\Big{\{}w\in\mathbb{P}(M\otimes\mathbb{C})\mid~{}\langle w,w\rangle=0,\langle w,\overline{w}\rangle>0\Big{\}}$ (5.1)
Let $\widetilde{\mathrm{O}}(M)=\ker(\mathrm{O}(M)\to\mathrm{O}(G_{M}))$ be the stable orthogonal group. It acts naturally on $\mathbf{D}(M)$ and the arithmetic quotient $\widetilde{\mathrm{O}}(M)\backslash\mathbf{D}$ is a normal quasi-projective variety, denoted by ${\rm Sh}(M)$. The tautological line bundle $\mathcal{O}(-1)$ on $\mathbf{D}(M)\subseteq\mathbb{P}(M\otimes\mathbb{C})$ is $\widetilde{\mathrm{O}}(M)$-invariant and descends to a line bundle on ${\rm Sh}(M)$, denoted by $\lambda_{M}$. It is called the Hodge line bundle on ${\rm Sh}(M)$.
@@
Set
$\Lambda_{g}=\left<2-2g\right>\oplus E_{8}(-1)^{\oplus 2}\oplus U^{\oplus 2}.$
Its discriminant group $\Lambda_{g}^{\vee}/\Lambda_{g}$ is a cyclic group of order $22g-$ 2. By global Torelli theorem for K3 surfaces, the period map induces an isomorphism
${\mathscr{F}}_{g}\cong{\rm Sh}(\Lambda_{g}).$
@@
which identifies the Noether-Lefschetz divisors on ${\mathscr{F}}_{g}$ as Heegner divisors on ${\rm Sh}(\Lambda_{g}).$
5.2 Heegner divisors on ${\rm Sh}(M)$
Let us consider the divisor class group of ${\rm Sh}(M)$.
@@ 5.2 Heegner divisors on ${\rm Sh}(M)$
For $m\in\mathbb{Q}$ and $\gamma\in G_{M}$, if $(m,\gamma)\neq(0,0)$, we define
$\mathrm{H}_{m,\gamma}=\widetilde{\mathrm{O}}(M)\backslash\sum\limits_{\begin{subarray}{c}v\in M+\gamma\\ \mathfrak{q}_{M}(v)=m\end{subarray}}v^{\perp},$
to be the Heegner divisor on ${\rm Sh}(M)$. If $(m,\gamma)=(0,0)$, then we may set $\mathrm{H}_{0,0}=-\lambda_{M}$.
In general, the Heegner divisor $\mathrm{H}_{m,\gamma}\in{\rm Cl}_{\mathbb{Q}}({\rm Sh}(M))$ is not irreducible. When $\gamma=-\gamma$ in $M^{\vee}/M$, the Heegner divisor $\mathrm{H}_{m,\gamma}$ has multiplicity two. The irreducible component of $\mathrm{H}_{m,\gamma}$ is the image of
$\Gamma_{v}\backslash\mathbf{D}(v^{\perp})\to\widetilde{\mathrm{O}}(M)\backslash\mathbf{D}$ (5.2)
for some $v\in M^{\vee}\subseteq M\otimes\mathbb{Q}$ satisfying $\mathfrak{q}_{M}(v)=m$ and $v\equiv\gamma\mod M$, where $\Gamma_{v}\subseteq\widetilde{\mathrm{O}}(M)$ is the stabilizer of $v$.
When $M$ contains two hyperbolic planes, by using Eichler’s lemma, there is a triangular relation between the Heegner divisors $\mathrm{H}_{m,\gamma}$ and its irreducible components. In this case, the span of all Heegner divisors $\mathrm{H}_{m,\gamma}$ equals the span of irreducible components of $\mathrm{H}_{m,\gamma}$.
5.3 Picard group of ${\rm Sh}(M)$
Let ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$ be the Picard group of ${\rm Sh}(M)$ with rational coefficients. As ${\rm Sh}(M)$ has only quotient singularity, ${\rm Sh}(M)$ is $\mathbb{Q}$-factorial and we have an isomorphism
${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))\cong{\rm Cl}_{\mathbb{Q}}({\rm Sh}(M)),$
which identifies $\mathrm{H}_{m,\gamma}$ as an element in ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$. Let us define
$\mathrm{ACusp}_{k}(\rho_{M})\subseteq{\mathrm{Mod}}_{k}(\rho_{M})$
as the subspace generated by elements in $\mathrm{Cusp}_{k}(\rho_{M})$ and the Siegel-Eisenstein series $E_{k,M}$, called the space of almost cusp forms of weight $k$ and type $\rho_{M}$. Then we have
Theorem 5.3 ([3]).
Assume that $n\geq 3$ and $M$ splits two hyperbolic planes. Then ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$ is spanned by the Heegner divisors and there is an isomorphism
$\Phi:{\rm Pic}_{\mathbb{C}}({\rm Sh}(M))\cong\mathrm{ACusp}_{\frac{2+n}{2}}(\rho_{M})^{\vee}$
by sending $\mathrm{H}_{m,\gamma}$ to the coefficient function
$f=\sum\limits_{\gamma\in G_{M}}\sum_{i\in\mathbb{Q}_{\geq 0}}c_{i,\gamma}q^{i}\mathfrak{e}_{\gamma}\mapsto c_{-m,\gamma},~{}\forall f\in\mathrm{ACusp}_{\frac{2+n}{2}}(\rho_{M}).$ (5.3)
As a consequence, we have
Theorem 5.4 (cf. [5, 6]).
The assumption on $M$ is the same as above. A finite linear combination $\mathrm{H}=\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}a_{m,\gamma}\mathrm{H}_{m,\gamma}$ is proportional to the Hodge line bundle $\lambda_{M}=\mathrm{H}_{0,0}$ if and only if
@@ 5.3 Picard group of ${\rm Sh}(M)$
for all cusp forms $\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}c_{m,\gamma}q^{m}\mathfrak{e}_{\gamma}\in\mathrm{Cusp}_{\frac{2+n}{2}}(\rho_{M})$.

### v2

@@ 5.1. Orthogonal Shimura variety and its compactification
Let $M$ be an even lattice of signature $(2,n)$. The Hermitian symmetric domain of type IV associated to $M$ is
@@
$\mathbf{D}^{\pm}=\left\{w\in\mathbb{P}(M\otimes\mathbb{C})\mid\langle w,w\rangle=0,\,\langle w,\overline{w}\rangle>0\right\},$
with two connected components. Fix a component $\mathbf{D}\subset\mathbf{D}^{\pm}$. We denote by $\mathrm{O}(M)(\mathbb{R})^{+}\leq\mathrm{O}(M)(\mathbb{R})$ the index-2 subgroup preserving $\mathbf{D}$. Given any arithmetic subgroup $\Gamma\leq\mathrm{O}(M)(\mathbb{R})^{+}$, we define the orthogonal Shimura variety
${\rm Sh}_{\Gamma}(M)\coloneqq\Gamma\backslash\mathbf{D}.$
Remark 5.1.
Note that $\Gamma^{\pm}\backslash\mathbf{D}^{\pm}\cong\Gamma\backslash\mathbf{D}$ for any $\Gamma^{\pm}\leq\mathrm{O}(M)(\mathbb{R})$ such that $\Gamma=\Gamma^{\pm}\cap\mathrm{O}(M)(\mathbb{R})^{+}$ has index 2 in $\Gamma^{\pm}$.
Let ${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))$ denote the Picard group of ${\rm Sh}_{\Gamma}(M)$ with rational coefficients. Since ${\rm Sh}_{\Gamma}(M)$ has only quotient singularities, it is $\mathbb{Q}$-factorial, and we have an isomorphism
${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))\cong{\rm Cl}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M)).$
The tautological line bundle $\mathcal{O}(-1)$ on $\mathbf{D}\subset\mathbb{P}(M\otimes\mathbb{C})$ descends to an element $\lambda_{M,\Gamma}$ in ${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))$, called the Hodge bundle. The Baily–Borel compactification $\overline{{\rm Sh}}_{\Gamma}(M)$ is constructed as
$\overline{{\rm Sh}}_{\Gamma}(M)={\rm Proj}\left(\bigoplus_{m\geq 0}\mathrm{H}^{0}({\rm Sh}_{\Gamma}(M),\lambda_{M,\Gamma}^{\otimes m})\right),$
endowing $\lambda_{M,\Gamma}$ with an ample extension $\overline{\lambda}_{M,\Gamma}$. Usually, we will simply write $\lambda_{M,\Gamma}$ if there is no further confusion.
Explicit Boundary Structure
The compactification admits a stratification:
$\overline{{\rm Sh}}_{\Gamma}(M)={\rm Sh}_{\Gamma}(M)\sqcup\bigsqcup_{I}\partial_{I}\sqcup\bigsqcup_{J}\partial_{J},$
where:
  • $I$ ranges over $\Gamma$-orbits of primitive isotropic lines in $M$
  • $J$ ranges over $\Gamma$-orbits of primitive isotropic planes in $M$
Each boundary divisor $\partial_{J}$ corresponds to a $\Gamma$-orbit of isotropic planes $J\subset M$, with the projection
$\pi_{J^{\perp}}:\mathbb{P}(M\otimes\mathbb{C})\setminus\mathbb{P}(J^{\perp}\otimes\mathbb{C})\to\mathbb{P}((M/J^{\perp})\otimes\mathbb{C})$
inducing the boundary map $\pi_{J^{\perp}}(\mathbf{D})\hookrightarrow\partial_{J}$.
In this paper, we primarily investigate two key cases that are closely related to moduli theory:
  1. (1)
    $\Gamma$ is a subgroup of the stable orthogonal group
    $\widetilde{\mathrm{O}}(M)\coloneqq\ker\left(\mathrm{O}(M)\to\mathrm{O}(G_{M})\right)\cap\mathrm{O}(M)(\mathbb{R})^{+}.$
    When $\Gamma=\widetilde{\mathrm{O}}(M)$, we write ${\rm Sh}_{\Gamma}(M)$ as ${\rm Sh}(M)$. Many moduli space of lattice-polarized hyper-Kähler varieties arises in this way.
  2. (2)
    $\Gamma=\Gamma_{0}:=\mathrm{O}(M)\cap\mathrm{O}(M)(\mathbb{R})^{+}$, the full orthogonal group. This case governs a lot of moduli spaces of bounded Calabi–Yau pairs $(X,D)$.
5.2. Heegner Divisors on ${\rm Sh}_{\Gamma}(M)$
We begin by recalling the construction of special divisors. For any vector $v\in M^{\vee}\subseteq M\otimes\mathbb{Q}$ with $\mathfrak{q}_{M}(v)<0$, the natural divisor $H_{v}$ on ${\rm Sh}_{\Gamma}(M)$ is defined as the image of the natural map
$\Gamma_{v}\backslash\mathbf{D}(v^{\perp})\hookrightarrow\widetilde{\mathrm{O}}(M)\backslash\mathbf{D},$ (5.1)
where $\Gamma_{v}\subseteq\widetilde{\mathrm{O}}(M)$ denotes the stabilizer subgroup of $v$.
Definition 5.2.
For a subgroup $\Gamma\leq\widetilde{\mathrm{O}}(M)$, $m\in\mathbb{Q}_{\leq 0}$, and $\gamma\in G_{M}$ with $(m,\gamma)\neq(0,0)$, the Heegner divisor is defined as
$\mathrm{H}_{m,\gamma}=\Gamma\Bigg{\backslash}\sum_{\begin{subarray}{c}v\in M+\gamma\\ \mathfrak{q}_{M}(v)=m\end{subarray}}v^{\perp}$
on ${\rm Sh}_{\Gamma}(M)$. Conventionally set $\mathrm{H}_{0,0}=-\lambda_{M,\Gamma}$ when $(m,\gamma)=(0,0)$. The subspace generated by Heegner divisors is denoted by ${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))^{\mathrm{Heegner}}$.
The Heegner divisor $\mathrm{H}_{m,\gamma}\in{\rm Cl}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))$ is generally reducible, and its irreducible components are natural divisors. In cases where $\gamma=-\gamma$ in $M^{\vee}/M$, the divisor $\mathrm{H}_{m,\gamma}$ acquires multiplicity two.
A fundamental property is the compatibility of Heegner divisors with pullback: for subgroups $\Gamma^{\prime}\leq\Gamma$, the natural projection $\pi:{\rm Sh}_{\Gamma^{\prime}}(M)\to{\rm Sh}_{\Gamma}(M)$ satisfies
$\pi^{*}\mathrm{H}_{m,\gamma}=\mathrm{H}_{m,\gamma}^{\Gamma^{\prime}},$
where $\mathrm{H}_{m,\gamma}^{\Gamma^{\prime}}$ denotes the Heegner divisor on ${\rm Sh}_{\Gamma^{\prime}}(M)$. Consequently, $\pi^{*}$ induces an isomorphism
${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))^{\mathrm{Heegner}}\cong{\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma^{\prime}}(M))^{\mathrm{Heegner}}.$
In particular, the space ${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))^{\mathrm{Heegner}}$ is independent of the choice of $\Gamma$.
@@ 5.2. Heegner Divisors on ${\rm Sh}_{\Gamma}(M)$
A crucial example is the Shimura varieties associated with the lattice
$\Lambda_{g}\coloneqq\langle 2-2g\rangle\oplus E_{8}(-1)^{\oplus 2}\oplus U^{\oplus 2},$
The global Torelli theorem for K3 surfaces provides a period map isomorphism
${\mathscr{F}}_{g}\cong{\rm Sh}(\Lambda_{g}),$
@@
establishing a correspondence between Noether–Lefschetz divisors on ${\mathscr{F}}_{g}$ and Heegner divisors on ${\rm Sh}(\Lambda_{g})$.
For a primitive vector $u\in M$, we call the positive generator $d$ of the ideal $\langle u,M\rangle$ the divisibility of $u$ and set $u^{*}\coloneqq u/d+M\in G_{M}$. A basic result is
Theorem 5.4 (cf. [BLMM17, Thm. 1.11]).
If $\dim{\rm Sh}_{\Gamma}(M)\geq 3$, then ${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))$ is spanned by natural divisors. Furthermore, if $M$ contains two hyperbolic planes and $\Gamma=\widetilde{\mathrm{O}}(M)$, then
${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))={\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))^{\rm Heegner}.$
Proof.
For the ease of readers, we recall the proof of the last assertion. When $M$ contains two hyperbolic planes, Eichler’s criterion says two primitive vectors $u$ and $v$ are lying in the same $\widetilde{\mathrm{O}}(M)$-orbit if $u^{2}=v^{2}$ and $u^{\ast}=v^{\ast}$. This implies the primitive Heegner divisor
$\mathrm{P}_{m,\gamma}:=\widetilde{\mathrm{O}}(M)\backslash\sum_{\begin{subarray}{c}v\in M\text{ primitive}\\ \mathfrak{q}_{M}(v)=m\\ v^{\ast}\equiv\gamma\pmod{M}\end{subarray}}\!\!\!v^{\perp}$ (5.2)
is irreducible. This shows that ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$ is generated by the primitive divisors $\mathrm{P}_{m,\gamma}$. As there is a triangular relation between the Heegner divisors $\mathrm{H}_{m,\gamma}$ and the primitive Heegner divisors, it follows that ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$ is spanned by $\mathrm{H}_{m,\gamma}$ as well.
When $M$ lacks two hyperbolic planes, Eichler’s criterion generally fails. It remains unknown whether ${\rm Pic}({\rm Sh}(M))$ is spanned by Heegner divisors. Nevertheless, if we consider the $\mathrm{O}(M)$-orbits, there is a similar result for $p$-elementary lattices.
Proposition 5.5 (generalization of Eichler’s criterion).
Let $p$ be a prime number. Let $M$ be a $p$-elementary even lattice of signature $(2,n)$ which satisfies one of the following conditions:
  • (i)
    $p\neq 2$ and $n>2$;
  • (ii)
    $p=2$ with $\ell(G_{M})\leq n$;
For any primitive vectors $u,v\in M$ satisfying
$u^{2}=v^{2}\quad\hbox{and}\quad u^{*}=v^{*}\in G_{M},$ (5.3)
there exists $g\in\Gamma_{0}$ such that $g(u)=v$.
Proof.
The case $p>2$ follows from [BC23, Thm. 5.15]. For $p=2$, we analyze two cases:
Case 1: $\ell(G_{M})\leq n-2$
In this case $M$ contains two hyperbolic planes. Eichler’s criterion directly applies since $U^{\oplus 2}\subset M$ guarantees transitivity on primitive vectors with matching invariants.
Case 2: $\ell(G_{M})=n$
We can decompose $M\cong U\oplus U(2)\oplus L(2)$ where $L=0$ or a negative definite unimodular lattice. Our argument depends on $L$’s parity. When $L$ is even, this can be proved by using Allcock’s trick (cf. [Allcock99, Lem. 1]). See also [KocaCaner2005OitA, Thm. 2.3.2] for a complete proof. If $L$ is odd, then there is a further decomposition (cf. [Ni79, Thm. 3.6.2])
$M\cong U\oplus U(2)\oplus A_{1}(-1)^{\oplus m}.$
We may fix the standard basis
  • $e_{1},f_{1}$ for $U$ with $\langle e_{1},f_{1}\rangle=1$,
  • $e_{2},f_{2}$ for $U(2)$ with $\langle e_{2},f_{2}\rangle=2$,
  • $\{r_{i}\}_{1\leq i\leq m}$ for $A_{1}(-1)^{\oplus m}$ with $r_{i}^{2}=-2$.
Let $u,v\in M$ be primitive with $u^{2}=v^{2}$ and $u^{*}=v^{*}\in G_{M}$. Let $d$ be their divisibility. Set $w=(u-v)/d\in M$ and make the decomposition:
$u=u_{0}+u_{R},\quad v=v_{0}+v_{R},\quad u_{0},v_{0}\in U\oplus U(2),\quad u_{R},v_{R}\in A_{1}(-1)^{\oplus m}.$
Then we are led to analyze the following subcases.
Subcase (i): $u_{0}^{*}\neq 0$, $v_{0}^{*}\neq 0$
As mentioned before, there exist $h,h^{\prime}\in\mathrm{O}(U\oplus U(2))$ such that $h(u_{0}),h^{\prime}(v_{0})\in U(2)\oplus A_{1}(-1)^{\oplus m}$. Extend to $M$ via block matrices:
$H=\begin{pmatrix}h&0\\ 0&I_{m}\end{pmatrix},\quad H^{\prime}=\begin{pmatrix}h^{\prime}&0\\ 0&I_{m}\end{pmatrix},$
where $I_{m}$ is the identity on $A_{1}(-1)^{\oplus m}$. Apply $H$ and $H^{\prime}$ to assume $u_{0},v_{0}\in U(2)$. Thus, $u,v\in U(2)\oplus A_{1}(-1)^{\oplus m}$.
Choose $u^{\prime},v^{\prime}\in U(2)\oplus A_{1}(-1)^{\oplus m}$ with $\langle u,u^{\prime}\rangle=d$, $\langle v,v^{\prime}\rangle=d$. Define Eichler transvections for isotropic $x$ and $y\perp x$:
$t(x,y)(\nu)=\nu-\langle y,\nu\rangle x+\langle x,\nu\rangle y-\frac{\langle x,\nu\rangle\langle y,y\rangle}{2}x,$
where $x,y\in M\otimes\mathbb{Q}$. The composition
$g=t(e_{1},-v^{\prime})\circ t(f_{1},w)\circ t(e_{1},u^{\prime})$
satisfies $g(u)=v$.
Subcase (ii): $u_{0}^{*}=0$, $v_{0}^{*}=0$
Similarly, assume $u_{0},v_{0}\in U$ via transformations as in (i). Thus, $u,v\in U\oplus A_{1}(-1)^{\oplus m}$. Take $u^{\prime},v^{\prime}\in U\oplus A_{1}(-1)^{\oplus m}$ with $\langle u,u^{\prime}\rangle=\langle v,v^{\prime}\rangle=d$. The composition
$g^{\prime}=t(e_{2},-v^{\prime})\circ t(f_{2},\frac{w}{2})\circ t(e_{2},u^{\prime})$
sends $u$ to $v$. The problem is that we have to show $t(f_{2},w/2)$ is an integral transformation. This requires:
    the divisibility of $w$ is even.
    $4\mid w^{2}$.
If $w$ does not satisfy these conditions, we claim that there exists $\phi\in\mathrm{O}(U\oplus A_{1}(-1)^{\oplus m})$ such that $(\phi(u)-v)/d$ satisfies the desired conditions. The transformation $\phi$ is achieved by possibly composing the following transformations:
  1. (1)
    The swap $\sigma:e_{1}\leftrightarrow f_{1}$.
  2. (2)
    For each $i$, the Eichler transvection $t(-e_{1},r_{i})$, which acts on basis as:
    $t(-e_{1},r_{i}):\begin{cases}e_{1}\mapsto e_{1}\\ f_{1}\mapsto f_{1}+e_{1}-r_{i}\\ r_{i}\mapsto r_{i}-2e_{1}\\ r_{j}\mapsto r_{j}\quad\forall j\neq i\end{cases}$
  3. (3)
    Sign changes $r_{i}\mapsto-r_{i}$ for indices with odd coefficients.
Here, we use $\sigma$ and $t(-e_{1},r_{i})$ to ensure that the divisibility of $w$ is even, while the sign change mapping ensures that $4\mid w^{2}$. Note that these transformations do not change the dual of these vectors in $G_{M}$.
Subcase (iii): $u_{0}^{*}=0$, $v_{0}^{*}\neq 0$
Here $d=2$. One can use the orthogonal transformations in (i) and (ii) to normalize our vectors without changing the norm and the dual in $G_{M}$. After normalization, assume:
$u_{0}=2e_{1}+u_{1},\quad u_{1}\in U(2),\quad u_{R}=r_{i}+\sum\limits_{j\neq i}c_{j}r_{j},\quad v_{0}\in U(2),$
for some $i$. As the transformation $t(-e_{1},r_{i})$ maps $(2e_{1}+r_{i})$ to $r_{i}$, it sends $u$ in $U(2)\oplus A_{1}(-1)^{\oplus m}$. Then proceed as in (i) to map $t(-e_{1},r_{i})(u)$ to $v$.
As an immediate consequence (together with Remark 5.1), we obtain the following.
Corollary 5.6.
Let $\pi:{\rm Sh}(M)\to{\rm Sh}_{\Gamma_{0}}(M)$ be the quotient map. The image of the pullback map
$\pi^{\ast}:{\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma_{0}}(M))\to{\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))$ (5.4)
is contained in ${\rm Pic}_{\mathbb{Q}}({\rm Sh}(M))^{\rm Heegner}$.
5.3. Heegner divisors as Coefficient Functionals
@@
Define the space of almost cusp forms of weight $k$ and type $\rho_{M}^{*}$
$\mathrm{ACusp}_{k}(\rho_{M}^{*})\subseteq{\mathrm{Mod}}_{k}(\rho_{M}^{*})$
as the subspace generated by cusp forms and the Siegel–Eisenstein series $E_{k,M}$.
Then we have
Theorem 5.8 ([Br14]).
Assume that $n\geq 3$ and $M$ contains $U\oplus U(N)$ as a direct summand. There is an isomorphism
$\Phi:{\rm Pic}_{\mathbb{C}}({\rm Sh}(M))^{\rm Heegner}\cong\mathrm{ACusp}_{\frac{2+n}{2}}(\rho_{M}^{*})^{\vee}$
realized by the coefficient functionals:
$\mathrm{H}_{m,\gamma}\longmapsto\Bigg{(}\underbrace{f=\sum_{\gamma^{\prime}\in G_{M}}\sum_{m^{\prime}}c_{-m^{\prime},\gamma}q^{m^{\prime}}\mathfrak{e}_{\gamma^{\prime}}}_{\text{Almost cusp form}}\ \mapsto\ \boxed{c_{-m,\gamma}}\Bigg{)}.$ (5.5)
Remark 5.9.
The assertion also holds for $M$ that does not split a hyperbolic plane. See [Br14, Thm. 1.4] and [Stein23, Thm. 1.1].
The geometric-modular correspondence established in Theorem 5.8 yields concrete criteria for detecting special divisors. A prime example is the characterization of the Hodge bundle through cusp form annihilation:
Theorem 5.10 (Hodge Bundle Criterion).
Under the assumptions of Theorem 5.8, a finite linear combination
$\mathcal{H}=\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}a_{m,\gamma}\mathrm{H}_{m,\gamma}$
is proportional to the Hodge line bundle $\mathrm{H}_{0,0}=-\lambda_{M}$ if and only if its associated coefficients functionals annihilate all cusp forms, i.e.
@@
for all cusp forms $\sum\limits_{\gamma\in G_{M}}\sum\limits_{m\in\mathbb{Q}}c_{m,\gamma}q^{m}\mathfrak{e}_{\gamma}\in\mathrm{Cusp}_{\frac{2+n}{2}}(\rho_{M}^{*})$.